# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ Abstract
Real-world time series data are inherently multivariate, often exhibiting complex inter-channel dependencies. Each channel is typically sampled at its own period and is prone to missing values due to various practical and operational constraints. These characteristics pose fundamental challenges related to channel dependency, sampling asynchrony, and missingness, all of which must be addressed to enable robust and reliable forecasting in practical settings. However, most existing architectures are built on oversimplified assumptions, such as identical sampling periods across channels and fully observed inputs at test time, which often do not hold in real-world scenarios. To bridge this gap, we propose ChannelTokenFormer, a Transformer-based forecasting model with a flexible architecture designed to explicitly capture cross-channel interactions, accommodate channel-wise asynchronous sampling, and effectively handle missing values. Extensive experiments on three benchmark datasets modified to reflect practical settings, along with one real-world industrial dataset, demonstrate the superior robustness and accuracy of ChannelTokenFormer under challenging real-world conditions.
@@ 1 Introduction
Accurate time series forecasting is critical in real-world applications such as industrial monitoring [17, 54], energy systems [46, 32, 14], and healthcare [12, 37], where predictive insights directly impact operational safety, resource efficiency, and long-term outcomes. This task is conventionally formulated under assumptions of fully observed data and equal sampling periods across channels. In contrast to these assumptions, time series data in some real-world applications are often far more complex than what existing models assume, as they typically consist of multivariate signals originating from heterogeneous sources such as sensors measuring temperature, pressure, or speed; control modules monitoring actuator positions or valve states; and biological systems recording heart rate, oxygen saturation, or neural activity. Owing to their diverse origins and functions, each of these channels may differ in sampling resolution and noise characteristics, while also being interdependent, reflecting intricate inter-channel dependencies, where the behavior of one channel can influence or correlate with others. Given these realities, the temporal dynamics of each channel can vary significantly, and modeling all channels in the same manner can be both unrealistic and potentially misleading—especially when inter-channel dependencies are simply added without addressing channel-wise heterogeneity. Yet many existing approaches [34, 42, 40, 4] presuppose idealized conditions—such as uniform temporal granularity across channels and fully observed inputs at test time—that seldom hold in practical deployments. Sensor systems in complex environments often consist of diverse devices, each governed by distinct physical constraints, design specifications, and application requirements. Different channels may therefore be sampled at varying periods; for example, one sensor might collect data every hour, while another records measurements every 15 minutes. Consequently, even when aligned over the same overall time window, such discrepancies lead to input sequences of varying lengths across channels, complicating both model design and preprocessing workflows.
@@
Figure 1: Newly proposed practical conditions include channel-wise asynchronous sampling periods and block-wise missing intervals at test time—further compound the long-standing challenge of inter-channel dependencies. Interpolation over coarsely sampled regions leads to signal distortion. See Appendix B for more details.
Despite these challenges, most standard time series forecasting models assume temporally aligned, fully observed inputs sampled at identical rates across channels. To conform to these assumptions, practitioners commonly resort to preprocessing techniques such as interpolation to impose regular intervals across channels, or downsampling, which reduces the resolution of high-frequency data to match the coarsest common sampling period across channels. While interpolation can introduce artificial continuity and spurious patterns, downsampling may cause loss of fine-grained temporal dynamics and eliminate informative fluctuations.
Beyond these structural discrepancies, real-world time series data are often further complicated by the presence of missing intervals at test time. In practice, sensors may malfunction temporarily or suffer communication failures, resulting in prolonged missing segments [48, 33], as opposed to sporadic individual gaps. Such contiguous absences pose a significant challenge in the input at test time, as accurate interpolation becomes difficult or even infeasible when missing intervals are long or the underlying signal is highly dynamic. In such cases, channel-dependent strategies offer a necessary alternative, enabling the model to infer missing dynamics from correlated signals across channels, rather than relying on unreliable interpolation.
@@
These challenges point to a clear need for forecasting models that remain robust under more realistic conditions as illustrated in Figure 1. In particular, models should be capable of: (i) processing multivariate time series with asynchronous sampling and variable-length inputs across channels, without requiring alignment or resampling; (ii) handling block-wise missing intervals during test-time inference without strictly relying on prior imputation; and (iii) leveraging inter-channel correlations through channel-dependent strategies to recover missing dynamics and preserve structural consistency.
To tackle these challenges, we propose ChannelTokenFormer, a Transformer-based forecasting framework designed to operate under partial observability and asynchronous sampling across channels. The key innovation is the use of channel tokens—summary representations that aggregate information from local temporal tokens within each channel. These tokens act as global attention anchors, facilitating both intra-channel modeling and selective cross-channel information exchange. Moreover, only the channel tokens are passed to the decoder, enabling prediction directly from compressed channel-wise representations—regardless of how many local tokens each channel contains. This design allows the model to flexibly accommodate variable-length local representations across channels. Rather than enforcing strict alignment or relying on interpolation, the model incorporates frequency-adaptive patching and masking-aware attention mechanisms to preserve the natural resolution of each channel and operate robustly even under block-wise missing inputs. In summary, our key contributions are as follows.
  • We identify three practical but often overlooked challenges in time series forecasting: inter-channel dependencies, channel-wise asynchronous sampling, and block-wise test-time missing intervals.
  • We propose ChannelTokenFormer, a Transformer-based model that introduces channel tokens to capture channel-specific dynamics. It supports frequency-adaptive patching to handle asynchronous and variable-length inputs across channels without alignment. Furthermore, it performs unified attention with masking strategies that enable structured information exchange between local and channel-wise global tokens, model inter-channel dependencies for robust forecasting, and address block-wise missing values at test time.
  • Extensive experiments on multiple realistic benchmarks and a real-world dataset demonstrate the model’s effectiveness under input asynchrony caused by differing sampling periods, test-time missing intervals, and inter-channel dependencies—hallmarks of real-world time series.
@@ 2 Related Work
Multivariate Time Series Forecasting: Channel-Dependent Strategy
Channel-Dependent (CD) strategy treats all channels jointly to capture inter-channel dependencies, and has been widely adopted across various neural architectures, including TCN/CNNs [30, 44], MLPs [11, 39, 16], Transformers [6, 25, 35, 42, 40, 53, 29, 45, 49], and GNNs [15, 47, 1, 27]. These methods are particularly effective when channels are strongly correlated, as they enable the model to exploit complementary information across channels. In practical settings, it is common to encounter channel-wise differences in sampling periods and block-wise missing intervals at test time. Under such conditions, CD strategies can be particularly beneficial, as they allow sparsely observed or low-resolution channels to benefit from denser or more informative ones through cross-channel interactions. In this sense, channel dependence is not only useful but may in fact be essential for robust forecasting under non-ideal conditions.
Irregularly Sampled vs. Channel-wise Asynchronous Time Series
Irregularly sampled multivariate time series refer to data in which each channel is observed at non-uniform and often unpredictable time points, resulting in irregular gaps within individual channels. Several recent works [22, 5, 51, 31, 26, 13, 36, 52, 18, 24, 20] address this challenge by developing models that operate directly on non-uniform time intervals. These approaches often incorporate time encoding schemes, interpolation-based alignment, attention mechanisms, or graph-based structures within forecasting or classification tasks. While these methods address general forms of irregular sampling, our focus is on a more structured and practical setting in which each channel is sampled at a fixed period, but the periods differ across channels. In our setting, each channel is sampled at regular intervals, though the sampling periods differ across channels. This reflects real-world sensor systems, where each channel is recorded at its own period depending on the task requirements or hardware constraints. As a result, the input sequence becomes naturally asynchronous, with sparse observations that follow a clear and structured pattern—reflecting fixed yet distinct sampling periods across channels, in contrast to the random and irregular missingness assumed in many previous studies.
@@
Early models such as BRITS [2] reconstruct missing values through imputation to reconstruct missing values based on bidirectional recurrent dynamics. General-purpose frameworks for time series, such as TimesNet [44] and TimeMixer++ [38], treat imputation as one of several downstream tasks. GRU-D [3] is designed for classification tasks on time series with missing values, while RAINDROP [52] extends this direction by explicitly modeling sensor-level missingness. However, in the context of forecasting tasks, the handling of test-time missing intervals—particularly block-wise gaps—remains underexplored. SERT [33] highlights the importance of handling block-wise missingness in real-world forecasting scenarios. While some studies [42, 56] evaluate model robustness under missing inputs, they typically replace missing values with zeros or mean values, without any explicit mechanism to address structured sparsity during test-time inference.
@@ 4 ChannelTokenFormer
Figure 2: Overview of ChannelTokenFormer (CTF). All tokens across channels pass through a unified attention layer, where local and global information is aggregated into channel tokens. Only the channel tokens are decoded by a decoder shared per sampling period to produce the final prediction.
Motivation for Channel Tokens
While some multivariate time series are regularly sampled, many real-world scenarios involve channel-specific sampling periods and prolonged test-time missing intervals. Such conditions limit the applicability of forecasting models that assume fixed-length, fully observed, and temporally aligned inputs. Although pre-filling is commonly used, it often distorts signal characteristics and induces artificial temporal patterns, ultimately degrading forecasting accuracy under asynchronous sampling and missingness.
To address these limitations, we propose Channel Tokens—learned global representations that summarize temporal dynamics within each channel and support forecasting under asynchronous sampling, test-time missingness, and inter-channel dependencies. Channel tokens attend to the observed patch-level inputs of their corresponding channels via masked self-attention, eliminating the need for aligned or interpolated sequences. This design enables the model to handle asynchronous sampling and missing intervals by compressing local information into channel tokens and enabling cross-channel interaction through attention among them.
@@
Each input channel is represented as a one-dimensional sequence $\mathbf{x}^{(i)}$. To determine an appropriate patch length, we apply the Fast Fourier Transform (FFT) to obtain the amplitude spectrum $\mathbf{A}^{(i)}=|\hat{\mathbf{x}}^{(i)}|$. For channel $i$, if a strong dominant frequency $f^{*}_{i}=\arg\max_{k}\mathbf{A}^{(i)}_{k}$ is detected, we set the patch length to its corresponding period: $\ell_{i}=\left\lfloor L_{i}/f^{*}_{i}\right\rfloor$. Otherwise, the patch length is determined based on the channel’s relative sampling period. This frequency-aware patching allows each channel to be divided into non-overlapping local patches that reflect its temporal characteristics, which are later summarized by channel tokens. Each local patch $\mathbf{p}^{(i)}_{j}\in\mathbb{R}^{\ell_{i}}$ is projected into a fixed-dimensional token $\mathbf{L}^{(i)}_{j}\in\mathbb{R}^{d}$ using a linear layer specific to its patch length: $\mathbf{L}^{(i)}_{j}=\mathbf{W}_{\text{patch}}^{(\ell_{i})}\mathbf{p}^{(i)}_{j}$. To improve efficiency, patches with the same length share a projection layer, enabling tokenizers to be reused across channels. Because patch counts and lengths vary across channels, this design balances efficiency and flexibility. In parallel, multiple channel tokens $\mathbf{C}^{(i)}=[\mathbf{C}^{(i)}_{1};\dots;\mathbf{C}^{(i)}_{C}]\in\mathbb{R}^{C\times d}$ are initialized for each channel to serve as global representations, where $C$ denotes the number of channel tokens for each channel. We add a fixed positional embedding $\mathbf{e}_{\text{pos}}$ and a learnable channel embedding $\mathbf{e}_{\text{ch}}^{(i)}$, which indicates the identity of channel $i$, to each local token, while the channel token receives only the channel embedding.
Input-Length Variability through Channel-wise Patch Masking
@@
Building on the flexibility of channel-wise dynamic patching, our framework allows for variable input lengths via patch masking. Leveraging this capability, we propose a patch masking strategy to handle interval missingness at test time. Specifically, if a missing segment spans an entire patch identified by the binary mask, the corresponding local token is excluded from the attention computation. As illustrated in Figure 2, this prevents fully unobserved patches from contributing to attention, thereby reducing the risk of propagating invalid information. Despite temporal gaps from masking, the model retains and understands temporal context through fixed positional embeddings. To improve robustness, we propose randomly masking a subset of channel-wise patches during training [28, 34]. This strategy improves resilience to test-time missing intervals and also serves as an implicit regularizer, reducing overfitting in low-resource forecasting scenarios.
Unified Attention via Masked Local and Channel Tokens

### v2

@@ Abstract
Real-world time series data are inherently multivariate, often exhibiting complex inter-channel dependencies. Each channel is typically sampled at its own period and is prone to missing values due to various practical and operational constraints. These characteristics pose three fundamental challenges involving channel dependency, sampling asynchrony, and missingness, all of which must be addressed simultaneously to enable robust and reliable forecasting in practical settings. However, existing architectures typically address only parts of these challenges in isolation and still rely on simplifying assumptions, leaving unresolved the combined challenges of asynchronous channel sampling, test-time missing blocks, and intricate inter-channel dependencies. To bridge this gap, we propose ChannelTokenFormer, a Transformer-based forecasting framework with a flexible architecture designed to explicitly capture cross-channel interactions, accommodate channel-wise asynchronous sampling, and effectively handle missing values. Extensive experiments on public benchmark datasets reflecting practical settings, along with one private real-world industrial dataset, demonstrate the superior robustness and accuracy of ChannelTokenFormer under challenging real-world conditions.
@@ 1 Introduction
Accurate time series forecasting is critical in domains such as industrial monitoring (Jean-Pierre et al., 2024; Zhao et al., 2024b), energy systems (Yao et al., 2025; Nascimento et al., 2023; Hu, 2024), and healthcare (He & Chiang, 2025; Tang et al., 2024), where predictive insights directly influence operational safety, resource efficiency, and long-term outcomes. Yet, real-world time series present several challenges that are often oversimplified by current modeling approaches.
One key challenge lies in the complex interdependencies among channels. Signals from different sensors or subsystems are rarely independent; their interactions often encode latent correlative dynamics crucial for accurate forecasting. Some studies have directly adopted channel-dependent architectures (Zhang & Yan, 2023; Wang et al., 2024c; Liu et al., 2024c), exploring inter-channel dependencies under specific structural assumptions. As an alternative, many models employ a channel-independent design (Nie et al., 2023; Wang et al., 2024b), which has shown competitive robustness to distributional drift (Han et al., 2024b). Nevertheless, carefully designed channel-dependent approaches (Chen et al., 2024a; Lee et al., 2024) can still exploit interdependencies to provide richer predictive signals and deliver performance gains beyond channel-independent baselines.
Another difficulty arises from the heterogeneity of data sources (Reiss et al., 2019; Filho et al., 2024; Dong et al., 2025; Ying et al., 2025; Agency, 2025). Time series signals commonly originate from diverse sensors, such as those tracking temperature, pressure, actuator positions, or biological signals. Due to differing physical properties and application contexts, these signals are frequently sampled at varying temporal resolutions, often leading to channel-wise (multi-source) asynchronous sampling in practice. Nonetheless, most existing approaches (Wang et al., 2024c; Chen et al., 2025) assume idealized input conditions: fully observed sequences sampled at identical intervals and aligned timestamps across channels. Such assumptions ignore variations in sampling periods and sequence lengths, complicating both model design and data preprocessing in real-world settings. Beyond channel-independent strategies, the challenge of channel-wise asynchronous sampling remains underexplored.
@@
Figure 1: Our proposed practical conditions highlight the simultaneous presence of channel-wise asynchronous sampling, block-wise missingness at test time, and inter-channel dependencies. Interpolation over coarsely sampled regions leads to signal distortion. See Appendix B for more details.
As with interdependencies and sampling variability, missing data represents another key challenge in time series modeling. Real-world signals often contain long contiguous intervals of missing observations, often arising from maintenance issues, sensor malfunctions, or communication failures (You et al., 2025; Nejad et al., 2024). In such cases, naive interpolation, which is commonly applied during preprocessing, can be unreliable or misleading, particularly for dynamic signals. More robust approaches can instead benefit from leveraging cross-channel correlations to infer missing dynamics, rather than relying solely on extrapolation from partially observed inputs.
@@
These three core challenges highlight the need for forecasting models that remain robust under realistic conditions, as illustrated in Figure 1. In particular, models should be capable of: (i) leveraging cross-channel dependencies to capture complex interactions and maintain structural consistency; (ii) handling asynchronous, variable-length inputs across channels without strict alignment or resampling; and (iii) addressing block-wise missing intervals during test-time inference by directly exploiting information from other channels, rather than relying on imputation methods.
While these requirements are clear, prior approaches have only provided partial solutions. Channel-dependent models capture inter-channel dependencies but overlook asynchrony and missingness; channel-independent strategies handle misalignment but lose cross-channel structure; and specialized approaches for missing values do not account for asynchrony, while irregular time series methods focus on within-channel non-uniformity under sparse settings rather than cross-channel heterogeneity in sampling.
To overcome these limitations, we propose ChannelTokenFormer, a unified Transformer-based forecasting framework designed to tackle all three challenges in a unified and simultaneous manner: channel-wise asynchronous sampling, test-time missing blocks, and cross-channel dependencies. The key is to revisit channel tokens under realistic multivariate conditions, viewing them as compact representations that aggregate local temporal information within each channel while also capturing cross-channel dependencies. Although this idea resembles the channel-level summary tokens in iTransformer (Liu et al., 2024c) and TimeXer (Wang et al., 2024e), our framework distinguishes itself by introducing a mask-guided attention strategy that enables a unified treatment of intra- and cross-channel interactions. In this way, the tokens serve as global attention anchors. Predictions are made from compressed channel-wise representations that account for heterogeneous sampling periods and masked tokens from missing blocks, regardless of the number of local tokens. This design preserves natural channel resolutions, remains robust to partially observed inputs, and adaptively captures cross-channel dependencies.
@@ 2 Related Work
Channel-Dependent Strategies for Multivariate Time Series
Channel-Dependent (CD) strategies jointly model all channels to capture inter-channel dependencies, and has been widely adopted across various neural architectures, including TCN/CNNs (Luo & Wang, 2024; Wu et al., 2023), MLPs (Han et al., 2024a; Huang et al., 2024), Transformers (Cheng et al., 2024; Liu et al., 2024b; Shu & Lampos, 2024; Wang et al., 2024e; c; Zhang & Yan, 2023; Liu et al., 2024c; Yang et al., 2024; Yu et al., 2023), and GNNs (Huang et al., 2023; Yi et al., 2023; Cao et al., 2020; Liu et al., 2022). Generative approaches such as COSCI-GAN (Seyfi et al., 2022) also employ CD principles, emphasizing cross-channel coherence in the context of sequence generation. These approaches are particularly effective when channels are correlated, since sparsely observed or low-resolution signals can benefit from denser ones through cross-channel interactions. In practice, however, most CD strategies only partially address real-world challenges: they typically assume aligned sampling across channels, are not designed to handle block-wise missing intervals and can be sensitive to noisy inter-dependencies. In contrast, our method utilizes channel tokens that explicitly encode cross-channel relationships while remaining robust under asynchronous sampling and block-wise missingness.
Irregularity of Multivariate Time Series (Multi-Source Asynchrony)
Irregularly sampled multivariate time series are characterized by channels observed at non-uniform and often unpredictable time points, resulting in irregular gaps within individual channels. A number of recent works (Che et al., 2018; Shukla & Marlin, 2021; Li et al., 2023a; Chen et al., 2023; Zhang et al., 2023; 2024; Mercatali et al., 2024; Liu et al., 2024a; 2025; Kim & Lee, 2024; Klötergens et al., 2025) address this challenge by developing models that operate directly on non-uniform time intervals, often using time encodings, interpolation-based alignment, attention mechanisms, or graph-based structures. Models such as Raindrop (Zhang et al., 2022) and Hi-Patch (Luo et al., 2025) are tailored for highly sparse settings, focusing on point-wise inter-channel relations rather than capturing temporal patterns. As a result, they are less suited to structured multi-source asynchrony and do not align with our proposed practical setting, which unifies three core challenges in continuous, long-term forecasting. In contrast, we target a real-world setting, including fixed but distinct sampling periods across channels, as categorized in the recent Time-IMM dataset (Chang et al., 2025).
@@
Models such as BRITS (Cao et al., 2018) and SAITS (Du et al., 2023) reconstruct missing values through imputation based on recurrent dynamics or self-attention, while general-purpose frameworks like TimesNet (Wu et al., 2023) and TimeMixer++ (Wang et al., 2024a) treat imputation as a downstream task. While these models can perform imputation, they typically rely on a separate forecasting module to complete the end-to-end pipeline, which increases complexity and may hinder practical deployment in real-world scenarios. BiTGraph (Chen et al., 2024b) and S4M (Jing et al., 2025) integrate missing-value handling into forecasting architectures, but still assume regularly sampled time series. Other studies such as TimeXer (Wang et al., 2024e) and TFT (Zhou, 2023) test robustness to missing inputs by simple replacements, offering no explicit mechanism for structured sparsity. SERT (Nejad et al., 2024) highlights the importance of handling block-wise missingness in real-world forecasting scenarios, reinforcing the need for explicit mechanisms beyond simple replacement. In contrast, our approach leverages channel tokens to minimize imputation-induced distortion and directly handle test-time missing intervals during inference, providing a robust solution for real-world settings.
@@ 4 ChannelTokenFormer
Repurposing Channel Tokens for Realistic Settings
Pioneering studies have explored the use of auxiliary tokens, for example, query tokens for extracting forecasting-relevant information from sequences of varying lengths (Kim et al., 2024), and special tokens for mediating information exchange across heterogeneous entities (Zhao et al., 2024a). Channel-level summary tokens have also been introduced in iTransformer (Liu et al., 2024c) and TimeXer (Wang et al., 2024e). Building on these insights, we repurpose channel tokens for realistic multivariate forecasting conditions, where patches are inherently unbalanced across channels due to heterogeneous sampling, optionally their own dominant frequencies, and prolonged missing intervals. In our framework, Channel Tokens act as compact abstractions that summarize each channel’s patches into stable channel-level embeddings, so that forecasting relies on these embeddings rather than on uneven patch sequences that would otherwise complicate the decoder input. This reframing allows the model to remain robust under practical conditions where patch imbalance poses a fundamental challenge.
@@
Prior frequency-guided patching (e.g., Moirai (Woo et al., 2024), LightGTS (Wang et al., 2025)) in time series foundation models targeted variable-length segmentation for univariate streams with heterogeneous granularities. In contrast, we repurpose this frequency-guided granularity selection for a channel-wise asynchronous multivariate setting. This enables non-aligned, per-channel patching while sharing projection layers (i.e., tokenizers) across equal patch lengths and introducing dedicated channel tokens, thereby preserving efficiency. Specifically, each input channel $\mathbf{x}^{(i)}$ is adaptively patched by estimating its dominant frequency via the Fast Fourier Transform (FFT). If a clear peak is identified, the patch length is set to the corresponding period; otherwise, it is chosen through a sampling-aware fallback. Full details on thresholds, dataset-specific examples, and the fallback procedure are provided in Appendix A.2. This frequency-aware patching allows each channel to be divided into non-overlapping local patches that reflect its temporal characteristics and are subsequently summarized by channel tokens. Each local patch $\mathbf{p}^{(i)}_{j}\in\mathbb{R}^{\ell_{i}}$ is projected into a fixed-dimensional token $\mathbf{L}^{(i)}_{j}\in\mathbb{R}^{d}$ using a linear layer parameterized by the patch length: $\mathbf{L}^{(i)}_{j}=\mathbf{W}_{\text{patch}}^{(\ell_{i})}\mathbf{p}^{(i)}_{j}$. In parallel, multiple channel tokens $\mathbf{C}^{(i)}=[\mathbf{C}^{(i)}_{1};\dots;\mathbf{C}^{(i)}_{C}]\in\mathbb{R}^{C\times d}$ are initialized per channel to serve as global representations, where $C$ denotes the number of channel tokens assigned to each channel. We add a fixed positional embedding $\mathbf{e}_{\text{pos}}$ and a learnable channel embedding $\mathbf{e}_{\text{ch}}^{(i)}$ that encodes channel $i$’s identity to each local token, while the channel token receives only the channel embedding.
Training-Time Proxying for Test-Time Missingness via Channel-wise Patch Masking
@@
Building on channel-wise dynamic patching, our framework supports variable input lengths via patch masking and explicitly targets block-wise missingness at test time. Unlike prior uses of random patch masking that primarily serve as a regularizer to curb overfitting, we repurpose it as a training-time proxy for realistic test-time incompleteness: inference-time inputs may be partially missing, so the model is trained to handle such missing blocks by design. We adopt random patch masking, as introduced in PatchDropout (Liu et al., 2023), applied to a subset of channel-wise patches at the input stage during training to simulate block-wise missingness at test time. As illustrated in Figure 2, if a missing segment spans an entire patch (as indicated by a binary mask) at test time, the corresponding local token is removed at the input stage, thereby being excluded from the attention computation. This prevents fully unobserved patches from contributing spurious signals, minimizing the risk of propagating invalid information. Despite temporal gaps induced by masking, temporal ordering and context are retained through fixed positional embeddings. This perspective shift improves resilience to test-time missing intervals and, as a side effect, acts as an implicit regularizer that mitigates overfitting in low-resource settings.
Figure 2: Overview of ChannelTokenFormer (CTF). All tokens across channels pass through a unified attention layer, where local and global information is aggregated into channel tokens. Only the channel tokens are decoded by decoders, each shared among channels with the same sampling period, to produce the final prediction.
A Unified Mask-Guided Attention with Channel Tokens for Real-World Challenges